OpenStudy (anonymous):

What is the standard form of the equation? 3y – 7 = –5x

OpenStudy (anonymous):

okay close it cause im here lol

OpenStudy (anonymous):

5x+3y=7

OpenStudy (anonymous):

Do you want me to make a new question or put it in the reply thing

OpenStudy (anonymous):

it don't matter. u can just put it in here if u want

OpenStudy (anonymous):

What is the slope of the line 4x – 8y – 16 = 0 in simplest form?

OpenStudy (anonymous):

m=1/5=.50

OpenStudy (anonymous):

1/2 sorry

OpenStudy (anonymous):

What is the y-intercept of the line represented by the equation? y = 7x + 2

OpenStudy (anonymous):

b=2

OpenStudy (anonymous):

how many more u got?

OpenStudy (anonymous):

5 more to go!

OpenStudy (anonymous):

wanna wait 5 min so i can take a break lol

OpenStudy (anonymous):

then ill get back on and finish em up with you

OpenStudy (anonymous):

K

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Ready?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Solve the system of equations. 2x+3y=1 y=3x+15

OpenStudy (anonymous):

(-4,3)

OpenStudy (anonymous):

hello?!

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

i was helping my mom

OpenStudy (anonymous):

@kylefoster58

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Mia deposits either $5 or $3.50 into her savings account. She made 14 deposits equaling a total of $59.50. How many times did she deposit $5 and how many times did she deposit $3.50? Mia made ? deposits of $5 and ? deposits of $3.50.

OpenStudy (anonymous):

Is it multiple choice?

OpenStudy (anonymous):

no sadly :*(

OpenStudy (anonymous):

okay. hang on

OpenStudy (anonymous):

Mia made 5 deposits of $5 and 7 deposits of $3.50

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

Which numbers belong to the range of the relation {(0, 7), (6, 0), (–3, –4), (–4, 7)}? Choose all answers that are correct.

OpenStudy (anonymous):

next?

OpenStudy (anonymous):

Which numbers belong to the range of the relation {(0, 7), (6, 0), (–3, –4), (–4, 7)}? Choose all answers that are correct.

OpenStudy (anonymous):

7,0,-4

OpenStudy (anonymous):

I KNEW IT!!!!!!

OpenStudy (anonymous):

Last one!

OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (1, –2) and has a slope of –3? y = ?x + ?

OpenStudy (anonymous):

oh crap!! i cannot do slope intercept haha i seriously have never been able to do it

OpenStudy (anonymous):

Gonna have to post that one as a new question bro. Sorry I couldn't get that one

OpenStudy (anonymous):

Its ok my friend i will let know know what i made

OpenStudy (anonymous):

just do me a favor and give me a medal lol

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

How do I?

OpenStudy (anonymous):

click best response

OpenStudy (anonymous):

I got a 66% :(

OpenStudy (anonymous):

nutuh forreal

OpenStudy (anonymous):

Off of ones I told you?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

what the heck! i used math way on the ones i want shure of and then the ones i knew i got i know i didn't get wrong

OpenStudy (anonymous):

Its ok Kyle at least you got a little over half right and having some right (although i failed ;( ) is better than having non right

OpenStudy (anonymous):

man thats bullcrap haha

OpenStudy (anonymous):

ik